5a) Explain whether each of the following relations on the set of real numbers is a function or not. For those (if any) that are indeed functions say whether they are one-to-one and/or onto. (2 marks)

i) y = f(x) = 2x2+1 xэR, y эR

ii) y = g(x) = 1/(x+1) (xэR, y эR , x != -1)

iii) Let h be a function from X = {1, 2, 3, 4} to Y = {a, b, c, d}.

h(1) = d, h(1) = c, h(2) = a, h(3) =b, and h(4) = b.

5b) Does either f or g have an inverse? If so, find this inverse. (1 marks)

5c) Find the composite functions f 。g and g。f . (2marks)

Question 6


1
Expert's answer
2021-08-30T16:42:46-0400

5a) Let us explain whether each of the following relations on the set of real numbers is a function or not. For those (if any) that are indeed functions let us say whether they are one-to-one and/or onto.


i) The relation "y = f(x) = 2x^2+1,\ x\in R, y \in R" is a function because of for each "x\in \R" there exists a unique "y\in\R" such that "y=2x^2+1=f(x)." Since "f(-1)=2(-1)^2+1=2\cdot 1^2+1=f(1)," this function is not one-to-one. Taking into account that the equation "0=2x^2+1" has no real solution, we conclude that this function is not onto "\R".


ii) The relation "y = g(x) = \frac{1}{x+1} (x\in \R, y \in\R , x \ne -1)" is a function because of for each "x\in \R\setminus\{-1\}" there exists a unique "y\in\R" such that "y= \frac{1}{x+1}=g(x)." Since "g(a)=g(b)" implies "\frac{1}{a+1}=\frac{1}{b+1}," and hence "a+1=b+1," we conclude that "a=b," and thus this function is not one-to-one. Taking into account that the equation "0=\frac{1}{x+1}" has no real solution, we conclude that this function is not onto "\R".


iii) Let "h" be a function from "X =\{1, 2, 3, 4\}" to "Y = \{a, b, c, d\}".

"h(1) = d,\ h(1) = c,\ h(2) = a,\ h(3) =b," and "h(4) = b."

This relation is not a function because of for "x=1" there exist two values "y=c" and "y=d" such that "y=h(x)."


5b) Since "f:\R\to \R" is not a bijection, it has no inverse function "f^{-1}:\R\to \R".

Since "g:\R\setminus\{-1\}\to \R" is not onto, it has no inverse function "g^{-1}:\R\to \R\setminus\{-1\}."

Taking into account that "f" is one-to-one and for any "y\ne 0" there exists a unique "x=\frac{1}{y}-1" such that "g(x)=\frac{1}{(\frac{1}{y}-1)+1}=y," we concvlude that there exist an inverse function "g^{-1}:\R\setminus\{0\}\to \R\setminus\{-1\}, \ g^{-1}(x)=\frac{1}{x}-1."



5c) Let us find the composite functions "f\circ g" and "g\circ f:"


"f\circ g(x)=f(g(x))=f(\frac{1}{x+1})=2(\frac{1}{x+1})^2+1=\frac{2}{(x+1)^2}+1;"


"g\circ f(x)=g(f(x))=g(2x^2+1)=\frac{1}{2x^2+1+1}=\frac{1}{2x^2+2}."